OpenStudy (anonymous):

For the given cost function C(x) = 28900+600 x + x^2 find: a) The cost at the production level 1850 b) The average cost at the production level 1850 c) The marginal cost at the production level 1850 d) The production level that will minimize the average cost e) The minimal average cost

OpenStudy (anonymous):

mostly i dont know how to solve d and e

OpenStudy (anonymous):

ANYONE?

OpenStudy (anonymous):

differentiate C(x) and equate this to zero solving this for x will give you the production level you require this is d)

OpenStudy (anonymous):

plugging this value into C(x) will give you (e)

OpenStudy (anonymous):

are you familiar with calculus ?

OpenStudy (anonymous):

YES I AM AND I TRIED TO DIFFERENTIATE C(X) AND THEN SET TO 0 AND SOLVE AND I DIDNT GET THE CORRECT ANS

OpenStudy (anonymous):

So what did you get when you took the derivative?

OpenStudy (anonymous):

ITS 600+2X FOR THE DERIVATIVE

OpenStudy (anonymous):

and with this derivative when u solve for x - it comes out to be -300. which is not correct according to my hw

OpenStudy (anonymous):

Hmm that seems as if it should be correct, are you sure you have the correct function?

OpenStudy (anonymous):

yes...its C(x) = 28900+600 x + x^2

OpenStudy (anonymous):

with the given function the derivative will be \[2x+600\] then set to 0 will give you \[x=-300 \]

OpenStudy (anonymous):

Average Cost = c' (x) / x = (2x + 600 ) / x = 2 + ( 600/ x )

OpenStudy (anonymous):

Chlorophyll how would i find x? set the average cost = 0?

OpenStudy (anonymous):

this gives -300 which isnt correct :(

OpenStudy (anonymous):

aah the average cost

OpenStudy (anonymous):

Hold on, I'm thinking!

OpenStudy (anonymous):

ok thanks a bunch!

OpenStudy (anonymous):

Average Cost = cost/ quantity

OpenStudy (anonymous):

http://faculty.wlc.edu/buelow/calc/nt4-7.html This is a site that explains this principle. It is easier to post this than to type it out.

OpenStudy (anonymous):

= ( 28,900+600 x + x^2 )/ x = x + 600 + ( 28,900 / x )

OpenStudy (anonymous):

@ sashley0915 thanks

OpenStudy (anonymous):

@sashley0915 That's just perfect :D

OpenStudy (anonymous):

Hope it helps... :)

OpenStudy (anonymous):

@rukh can you continue?

OpenStudy (anonymous):

im trying

OpenStudy (anonymous):

sigh....

OpenStudy (anonymous):

Okie, let me give more hints ;)

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

Take derivative of average cost Avg (x) = x + 600 + ( 28,900 / x ) Avg' (x) = ( - 28,900 / x² ) + 1 = 0 => x = √ 28,900 = 170 units

OpenStudy (anonymous):

Now are you fine now?

OpenStudy (dumbcow):

\[\text Avg cost = \frac{C(x)}{x}\] differentiate using quotient rule, and set equal to zero \[\frac{x*C'(x) - C(x)}{x^{2}} = 0\] \[x*C'(x) - C(x)= 0\] \[C'(x) = \frac{C(x)}{x}\] so you see that avg cost is minimized when it equals marginal cost

OpenStudy (anonymous):

thanks everyone